### v2

# Vascular Ageing and Smoking Habit Prediction via a Low-Cost Single-Lead ECG Module

Abstract
This paper presents a novel low-cost method to predict: i) the vascular age of an apparently healthy young person, ii) whether or not a person smokes, using only the lead-I of the electrocardiogram (ECG). We begin by collecting (lead-I) ECG data from 42 subjects (male, female, smoker, non-smoker) aged 18 to 30 years, using our custom-built low-cost single-lead ECG module, and anthropometric data, e.g., body mass index, smoking status, blood pressure etc. Under our proposed method, we first pre-process our dataset by denoising the ECG traces, followed by baseline drift removal, followed by z-score normalization. Next, we divide ECG traces into overlapping segments of five-second duration, which leads to a 145-fold increase in the size of the dataset. We then feed our dataset to a number of machine learning models, a 1D convolutional neural network, a multi-layer perceptron (MLP), and ResNet18 transfer learning model. For vascular ageing prediction, Random Forest method outperforms all other methods with an $R^{2}$ score of 0.99, and mean squared error of 0.07. For the binary classification problem (smoker or non-smoker), XGBoost method stands out with an accuracy of 96.5%. Finally, for the 4-class classification problem that aims to differentiate between male smoker, female smoker, male non-smoker, and female non-smoker, MLP method achieves the best accuracy of 97.5%. This work is aligned with the sustainable development goals of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged.
Index Terms:
Electrocardiogram, vascular age, single-lead ECG, smoking, low-cost sensors, healthcare sensing.
I Introduction
Vascular age (also known as heart age or ECG-age) refers to the natural deterioration in structure and function of blood vessels as a person gets older [1]. In other words, as blood vessels age, they become stiffer, less elastic, and more prone to plaque buildup. Vascular ageing results from factors such as ageing, high blood pressure, high cholesterol, and smoking. Vascular ageing is a major risk factor for a number of cardiovascular diseases (CVD) and other diseases, including the following: hypertension (high blood pressure), atherosclerosis (narrowing and hardening of arteries), stroke, heart attacks, heart failure, endothelial dysfunction (impaired function of the inner lining of blood vessels), organ damage (due to reduced blood supply to organs, e.g., kidneys, eyes, etc.), and cognitive decline (due to reduced blood flow to the brain) [2], [3].
Among various risk factors, smoking is one significant factor that impacts the vascular age, and thus, impacts the cardiovascular health. Chronic smoking could lead to various alterations in the electrocardiogram (ECG), e.g., increased heart rate (tachycardia), increased risk of developing cardiac arrhythmias (such as premature atrial contractions and premature ventricular contractions), changes in the QT interval and ST-segment on the ECG (which may indicate myocardial ischemia, i.e., insufficient blood flow to the heart muscle), increased risk of coronary artery disease (shown by ST-segment depression and T-wave inversion), vasoconstriction (narrowing of blood vessels), and reduced oxygen levels [4].
As discussed above, vascular age and smoking both result in alterations in the ECG [4], [5]. This motivates us to study the feasibility of using ECG as a clinical tool to assess the impact of vascular age and smoking on cardiovascular health. This objective is inline with the broad research efforts that aim to use ECG as a diagnostic tool with the ultimate objective of bringing down the annual CVD-based deaths by early detection of CVDs [6]. A reduction in proliferation of CVDs will imply a reduction in health burden of countries in lieu of health insurance costs. However, note that manual assessment of the ECG to look for possible irregularities in the heart rhythm due to vascular ageing and smoking is time-consuming and error-prone. Hence, artificial intelligence (AI)-based methods which could automatically and reliably predict the vascular age [7] and smoking status of a person (smoker or non-smoker) from the (single-lead) ECG data are the need of the hour. This is precisely the scope of this work.
Existing methods for vascular ageing prediction and the dilemma: Vascular age prediction has traditionally been done using invasive methods which use catheters and pressure sensors placed directly inside arteries to obtain arterial stiffness and central aortic blood pressure measurements. Recently, there has been growing interest in designing non-invasive methods, which utilize either photoplethysmography (PPG) or ECG signals for vascular ageing prediction, [8], [7]. Among them, pulse wave velocity (PWV) method and its variants are considered the gold standard for measuring arterial stiffness. PWV method measures the speed at which the pressure wave travels along the arterial tree between two points. One important variant is carotid-femoral PWV (cfPWV) method which reflects the stiffness of the central arteries. Brachial-Ankle PWV (baPWV) method is another important variant that measures the time taken for the pressure wave to travel between the brachial and ankle arteries. Pulse wave analysis (PWA) method is another popular method that analyzes the shape of the arterial pulse wave to derive indices related to arterial stiffness and vascular function. Other than that, there are specific methods that examine the endothelial function, e.g., ultrasound-based flow-mediated dilation method, peripheral arterial tonometry to obtain reactive Hyperaemia index, etc. However, each of these methods has its drawbacks, e.g., invasive, high cost, need for trained professionals and/or specialized equipment, inconsistent results, etc.
Contributions.
  • We design a low-cost single-lead ECG sensor module (with a price tag of 10 USD), and utilize it to collect (lead-I) ECG data from 42 subjects (male, female, smoker, non-smoker) aged 18 to 30 years.
  • We feed our custom dataset to various machine, deep and transfer learning methods, which reliably predict the vascular age of a person using a low-cost single-lead (lead-I) ECG sensor module only. This is a giant leap compared to the related work that either utilizes the 12-lead ECG [5] or utilizes invasive methods (e.g., central aortic blood pressure monitoring) for the same task.
  • We also pass our dataset to a number of machine and deep learning methods, which reliably differentiate between a smoker and a non-smoker solely based on the single-lead ECG data. To the best of our knowledge, this work is the first that aims to infer the smoking status of a person from his/her single-lead ECG data.
Finally, note that this work is also aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged populations (e.g., people in developing countries and people in remote areas).
Outline. Section II provides a summary of the selected related work. Section III describes the proposed methodology as well as the hardware architecture of our custom low-cost single-lead ECG module. Section IV discusses the data collection process for our custom ECG dataset, the data pre-processing steps, and feature extraction process. Section V enumerates the ML, DL and TL models we have implemented. Section VI describes selected results for both problems (vascular ageing prediction and smoking habit prediction). Finally, Section VII concludes the paper.
II Related work
To date, a multitude of methods have been investigated, which either utilize the 12-lead ECG [5] or PPG sensors [9].
Statistical methods: [10] develops a Bayesian statistical model that predicts an individual’s heart age based on his/her resting 5 min 12-lead ECG. They evaluated their model on 776 healthy individuals, 221 with cardiovascular risk factors, 441 with cardiac disease, and a small group of highly endurance-trained athletes. Their model predicts a higher heart age for the subjects with risk factors, with heart diseases, and athletes. Motivated by the situations where cfPWV cannot be measured, [11] computes the estimated pulse wave velocity (ePWV) from the age and brachial blood pressure. They measure cfPWV in two-hundred and fifty-two adults, and compare it with each participant’s ePWV, in order to correlate the ePWV with cfPWV (which is an established measure of vascular aging).
Machine learning based methods: Recently, machine learning methods for vascular ageing prediction have gained widespread attention (see the survey article [7] and references therein for a quick review of the development in the field). Due to space constraints, we discuss selected works below. [12] utilizes a 13-years MESA dataset of 6459 atherosclerotic CVD-free participants and demonstrates that their proposed support vector machine (SVM)-based CVD risk calculator outperforms the classical PCE-based CVD risk calculator. [13] utilizes a public dataset of video PPG signals acquired from smartphones, does the feature extraction, and trains a number of machine and deep learning methods in order to predict healthy vascular ageing. Authors of [14] take a critical approach and study the bias problem in ML-based vascular age prediction methods. They argue that overfitting and lack of large datasets (with diverse ethnic populations) might be the reasons behind the fact that ML methods yield inaccurate results during the clinical trials. Similarly, [15] argues that it is possible to accurately predict patients’ Bayesian 5-min ECG Heart ages from their standard, resting 10-s 12-lead ECGs using standard regression models only, without relying upon the AI techniques. They utilize a public dataset with a total of 2,771 subjects (1682 healthy volunteers, 305 with cardiovascular risk factors, 784 with cardiovascular disease).
Deep learning based methods: In [16], authors train a residual neural network on the 12-lead ECG data from the CODE study cohort (1,558,415 patients), and report that the patients with a heart age gap of more than 8 years have a higher mortality rate. [17] trains a convolutional neural network (CNN) on 12-lead ECG data of 17,042 subjects, and reports that the age predicted by their model matches well with the chronological age of young population, but not for the elderly (with age 6060$\ge$ 6060 years). [18] trains the ECG12Net model (basically, a CNN with 82 convolutional layers) on 71,741 subjects (20 to 80 years old), and reports that the high ECG-age is not only correlated with mortality but with several CVD-related outcomes as well. Authors of [19] collect PPG data from 757 participants, utilize a Gaussian mixture model to decompose the PPG signal into incident and reflected waves for feature extraction, and eventually train a shallow artificial neural network for vascular ageing prediction. [20] trains a CNN on 10-second 12-lead ECG data and CVD-related metadata from 499,727 subjects, with the aim to learn the genetic basis of cardiovascular ageing, in order to support future personalised/precision medicine (when genomic information is readily available). In [21], researchers train their CNNs using 10-second samples of 12-lead ECG signals from 499,727 patients in order to predict their biological age and gender. Authors of [22] train a CNN on over 700,000 ECGs from the Mayo Clinic, U.S.A, in order to later predict the ECG-age of 4,542 participants in two cities in Russia. Authors report that $\delta$-age (i.e., ECG-derived artificial intelligence-estimated age minus chronological age) biomarker is strongly correlated with the CVD risk factors, e.g., blood pressure, BMI, total cholesterol and smoking. [23] trains a CNN on the 12-lead ECG data of 774,783 unique subjects, and conclude that the patients with peripheral microvascular endothelial dysfunction (an index of vascular aging) and accelerated physiologic aging are at a greater risk for CVDs.
Classical CVD risk calculator methods: In addition to vascular ageing prediction methods, other methods with related objectives also exist. For example, the Framingham risk score (FRS) [24], pooled cohort equations (PCE) [25], and systematic vascular risk evaluation (SCORE) are few widely used tools that take into account various risk factors such as gender, systolic blood pressure, smoking habits, and HDL cholesterol in order to estimate an individual’s 10-year risk of developing a cardiovascular event, such as coronary heart disease or stroke. However, it has been reported that the FRS method overestimates the risk in young age groups, while the PCE method overestimates the risk in older age groups [26].
Low-cost single-lead ECG module design: Since this work utilizes the AD8232 chip (a low-cost single-lead ECG sensor with a price tag of 5 USD), it is imperative to summarize the works that design low-cost single-lead ECG module with the unanimous aim to provide affordable sensing solutions for remote health monitoring. In [27],[28], authors consider a AD8232 chip-based low-cost single-lead ECG module for atrial fibrillation detection (a common cardiac arrhythmia associated with stroke risk). [29] presents a cost-effective method for detecting driver drowsiness, utilizing heart rate variability data from the ECG sensor. [30], [31], [32] all focus on the design of an internet-of-things (IoT)-based low-cost ECG system which reports various biomarkers (e.g., pulse rate, SpO2, body temperature, etc) of remote Covid-19 patients to the doctor, and could detect various cardiac abnormalities such as premature ventricular contraction, premature atrial complex, etc. [33] aims to estimate blood glucose level in a subject using a non-invasive wearable ECG device, using features such as T-wave, ST-segment, etc.
III The Methodology and the Measurement Equipment
III-A The Proposed Framework
Fig. 1 provides a detailed overview of our proposed method. As can be seen, we utilize our custom low-cost single-lead ECG sensor module in order to construct a labelled dataset. We then do pre-processing of the noisy ECG data, followed by segmentation, followed by feature extraction. Eventually, we pass a total of 21 features (13 ECG-based features and 8 other features) to a number of ML algorithms, a 1D CNN model, and the ResNet18 model which do vascular ageing prediction and smoking habit prediction.
Figure 1: Our proposed method for vascular ageing and smoking habit prediction.
III-B The Low-cost Single-Lead ECG Module
We utilize the AD8232 chip which is a low-cost single-lead ECG sensor (with a price tag of 5 USD). This ECG sensor requires a DC voltage in the range of 2-3.5 volts, and comes with three electrodes which enable it to capture the tiny voltages generated by the human heart, at a rate of 100 Hz. For systematic data acquisition, the AD8232 ECG sensor is interfaced with the ESP32-WROOM micro-controller board (developer version). Further, a 3.7 volts lithium-ion battery is used to power the ESP32 board, the ECG sensor, and TP4056 module (basically, a battery management system). This way, we realize a very cost-effective single-lead ECG sensor module (with a price tag of 10 USD). We connect and solder all the aforementioned hardware components on a veroboard, and then place it in an acrylic body designed via the Solidworks tool. We program the ESP32-WROOM and AD8232 ECG sensor using the Arduino IDE with the objective to read and log the ECG waveform for the purpose of our custom dataset construction. Fig. 2 shows the complete schematic of our custom low-cost single-lead sensor ECG Module.
Figure 2: Our custom single-lead ECG sensor module.
IV Data Collection, Data Pre-Processing and Feature Extraction
IV-A Data Collection
Fig. 3 (left subplot) shows the theoretical configuration of a three-electrode single-lead ECG sensor module, that forms the einthoven triangle for the purpose of acquiring the lead I of the ECG signal. Fig. 3 (right subplot) shows a screenshot from our data collection campaign where it can be seen that the positive (red) electrode was placed on the right wrist, the negative (yellow) electrode was placed on the left wrist. Finally, the ground (green) electrode was placed on the right leg (though it is not visible in Fig. 3).
Figure 3: The three electrodes are placed such that together they form the einthoven triangle to acquire the lead I of the ECG (left subplot). The right subplot shows a snippet from our data collection campaign.
We utilized our custom module to collect single-lead ECG data from 42 subjects (26 male, 16 female) of age 18 years to 30 years, on campus [^1] This work was approved by the ethical institutional review board (EIRB) of Information Technology University (ITU), Lahore, Pakistan.. Each subject contributed to 3 minutes worth of data, leading to a dataset having 126 minutes (7560 seconds) of single-lead ECG data. Among the male subjects, 16 were smokers, while 10 were non-smokers. Similarly, 4 female subjects were smokers, while 12 were non-smokers. For the ground truth for the vascular ageing prediction problem, we utilized chronological age as a proxy for the vascular age (which is a reasonable assumption keeping in mind that we collected data from apparently healthy young subjects only) [15],[13]. In addition to acquiring the ECG traces of the subjects, we also recorded the relevant metadata such as age, height, weight, heart disease history in the family, smoking habits, sleep duration, blood pressure (systolic and diastolic), body mass index (BMI), and the resting heart rate.
IV-B Data Pre-Processing
We first implemented a Butterworth low-pass filter with a cutoff frequency of 18 Hz and order 3, to denoise the ECG traces. We then removed the baseline wander (via a median filter), followed by the z-score normalization of the clean ECG traces. The onset of the P wave was used to detect the start of an ECG cycle (i.e., the PQRST wave). Anomalies in the ECG traces (that arise due to motion artifacts and due to improper placement of electrodes and cause a streak of zeros, not a number (NaN) values, or abnormal patterns) were manually identified and removed, from the onset of the P wave in that PQRST interval to the onset of the P wave in the next interval.
Segmentation: Each of the 42 examples/recordings were further divided into a number of five-second long (overlapping) segments with a stride of one second. This led to an increase in number of examples from 42 to 6131. Table I outlines the distribution of our dataset after segmentation.
Table I: Distribution of our dataset (after segmentation)
Class Label No. of Examples
Male $3455$
Female $2676$
Total $6131$
Male Smokers $2111$
Female Smokers $734$
Total Smokers $2845$
Male Non-Smokers $1344$
Female Non-Smokers $1942$
Total Non-Smokers $3286$
IV-C Feature Extraction
It is well-known that the ECG signal (with P,Q,R,S,T points marked) contains a wealth of information that has clinical significance to diagnose a number of heart abnormalities, including the condition of the heart, i.e., vascular ageing. Thus, the feature extraction from ECG involves the detection of P wave, QRS complex, T wave, as well as the measurement of the duration of a number of intervals between the marked points, e.g., RR interval (see Fig. 4 that shows one typical cardiac cycle with various intervals annotated).
Figure 4: One cardiac cycle of a typical ECG signal annotated with many intervals (features) that have clinical significance.
We extracted a number of features, e.g., RR interval, QT interval, P wave duration, PP interval, TP interval, PR interval, T wave duration, ST segment, QRS complex duration, PR segment, etc., from the segmented ECG data (using the Neurokit2 library). The extraction process involved applying a discrete wavelet transform to the ECG signals, which helped focus on the various peaks in the PQRST wave (i.e., one cardiac cycle) and the corresponding intervals.
Computing the various intervals allowed us to compute a few derived features as well. For example, we utilized the QT interval and RR interval to compute the corrected QT interval (QTc) as follows: $QTc=\frac{QT}{\sqrt{RR}}$. Further, using the RR interval, we derived the root mean square of successive differences (RMSSD) (which measures the variability between adjacent RR intervals), and the standard deviation of RR intervals (SDNN) (which quantifies the dispersion of the RR intervals).
Recall that each ECG segment is of duration 5 seconds, thus, each segment typically consists of about 5-10 cardiac cycles. Therefore, though we initially did feature extraction on a cardiac-cycle basis, we then averaged all the 13 extracted features across a segment in order to achieve reliable and stable features. We further observe that the values of the extracted features lie within the nominal range specified by the American heart association for the healthy young people [34]. Finally, Table II presents the values of the extracted features for one example from our dataset (i.e., corresponding to one segment of a male subject), for illustration purposes.
ECG Features Value (ms)
RR Interval $712.32$
RMSSD $65.59$
SDNN $52.40$
QT Interval $362.33$
P-wave Duration $101.47$
PP Interval $777.88$
TP Interval $211.74$
PR Interval $135.79$
T-wave Duration $176.13$
ST Segment $59.36$
QRS Complex Duration $126.84$
Corrected QT Interval $429.31$
PR Segment $34.32$
Table II: An illustration of the feature extraction process (i.e., one example from our dataset, corresponding to one segment of a male subject).
V Machine, Deep and Transfer Learning Models Implemented
Machine learning models: For the vascular ageing prediction problem, we implemented the following ML models: linear regression, ridge regression (linear regression with regularization), decision tree regression, and random forest regression (an ensemble of decision trees). For smoking prediction problem, we implemented the following ML models: decision tree, random forest, Gaussian naive bayes, and XGBoost.
Deep learning models: For the vascular ageing prediction problem, we implemented a 1D CNN (using the Keras library). The model utilizes 32 filters and a kernel size of 3, followed by a ReLU activation function. In the fully connected layers, a ”Dense” layer with 128 neurons and a ReLU activation are employed, followed by a dropout layer with a rate of 0.2 to mitigate overfitting. Lastly, in the output layer, a ”Dense” layer with a single neuron and a linear activation function is used for regression. The model is compiled using the Adam optimizer, with 220 epochs and a batch size of 23. For smoking prediction problem, we implemented a multi-layer perceptron (MLP) with one hidden layer having 10 neurons. We used the tanh activation function, and set the regularization parameter Alpha to 0.01. (We did feature selection and passed 8 most significant features to the MLP).
Transfer learning model: For the vascular ageing prediction problem, we did transfer learning by utilizing ResNet18, a popular CNN model in computer vision community. The model has 18 layers, and takes as input the RGB images of size 224x224 pixels. The structure of ResNet-18 is as follows: i) convolutional layer (7x7, stride 2), for feature extraction; ii) max-pooling (3x3, stride 2), to reduce the spatial dimensions of the feature maps; iii) residual blocks (5 groups with 2 or 3 basic blocks each) with skip connections [^2] Each BasicBlock layer comprises a convolutional layer, followed by batch normalization, followed by a ReLU activation function.; iv) global average pooling, to further reduce the spatial dimensions of the feature maps; v) a fully connected layer with 512 input features, followed by a ReLU activation, an output layer with a single neuron to predict the vascular age of the input sample.
Training of ML, DL and TL models: The dataset (consisting of 6131 samples) was divided into three parts. 60% (i.e., 3679 samples) of the dataset was used for training, 20% (i.e., 1226 samples) for validation, and remaining 20% (i.e., 1226 samples) for testing. For each method, the best combination of hyperparameters was identified by conducting a grid search. For transfer learning, we translated the 1D ECG segments into spectrograms and fed them to the ResNet18 model.
VI Results
VI-A Performance Metrics
For the regression problem (of vascular aging prediction), we utilize the mean squared error (MSE) and $R^{2}$ score as performance metrics. Note that a higher $R^{2}$ score means a better fit and a lower MSE means a better model predictive accuracy. For the classification problem (of smoking habit prediction), we construct confusion matrices which provide details such as true positives, true negatives, false positives, and false negatives.
VI-B Vascular Ageing Prediction
We begin by doing the feature correlation analysis. Fig. 5 reveals that vascular ageing is agnostic to the following features: BMI, sleep duration, RMSSD, SDNN and PR interval. Similarly, we learn from Fig. 5 that following features are most relevant to the vascular ageing: heart rate, P-wave duration, QT interval, PP interval, and T-wave duration.
Figure 5: Feature correlation analysis (target variable is age).
Among all the ML models, the linear models, i.e., linear regression model and its variant ridge regression performed the worst with an MSE of 5.79, $R^{2}$ score of 0.337, and MSE of 5.24, $R^{2}$ score of 0.37, respectively. This hints at a possible non-linear relationship between the ECG-based features and the vascular age. This hypothesis is later verified as the decision tree model and random forest model (an ensemble of decision trees) performed very well with a very small MSE of 0.258, high $R^{2}$ score of 0.97, and small MSE of 0.07, high $R^{2}$ score of 0.99, respectively. On the deep learning front, the 1D CNN also performed well with an MSE of 0.251 and an $R^{2}$ score of 0.971 (at a cost of more computational complexity). Finally, ResNet18 model achieved an MSE of 1.191, and a $R^{2}$ score of 0.86. Table III (bottom half) summarizes the MSE and $R^{2}$ results for all 6 of our models.
Table III also highlights that our ML models (random forests and decision tree) and DL model (1D CNN) provide superior MSE and $R^{2}$ performance compared to the 12-lead ECG-based state-of-the-art methods for vascular aging prediction.
Authors Year Waveform Model implemented Age Range Performance Metric
Starc et al. [35]
2012
12-lead ECG
Multiple Linear Regression
3-100
R2=0.76
Lima et al. [16]
2021
12-lead ECG
Residual Network
20-80
MAE=8.83
Chang et al. [18]
2022
12 Lead ECG
Deep Learning
20-80
MAE=6.899 R2=0.822
Zvuloni et al. [36]
2023
12-lead ECG
FE and DL
20-85
MAEFE=10.6 R2FE=0.60 MAEDL=6.32 R2DL=0.83
Ours 2023 1-lead ECG Random Forest 18-30 MSE=0.07; R2=0.99
Ours
2023
1-lead ECG
Decision Tree
18-30
MSE=0.258; R2=0.97
Ours
2023
1-lead ECG
Ridge Regression
18-30
MSE=5.24; R2=0.37
Ours
2023
1-lead ECG
Linear Regression
18-30
MSE=5.79; R2=0.337
Ours
2023
1-lead ECG
1D CNN
18-30
MSE=0.251; R2=0.971
Ours
2023
1-lead ECG
ResNet18
18-30
MSE=1.191; R2=0.86
Table III: MSE and $R^{2}$ accuracy of our ML, DL and TL models and their comparison with the previous work.
Fig. 6 plots the histograms of the error in vascular aging prediction for all 6 models we have implemented. We observe that the random forest, the decision tree, and 1D CNN have the smallest error spread among all, as expected.
Figure 6: Error Distribution of our ML, DL and TL models.
VI-C Smoking Habits Prediction
We again do feature correlation analysis first. Fig. 7 reveals that smoking makes the least impact on the following features: QRS complex duration, PR interval, QTc, and heart disease history in the family. Similarly, we learn from Fig. 7 that following features alter the most due to smoking: gender, ST segment, RR interval, systolic blood pressure, and heart rate.
Figure 7: Feature correlation analysis (target variable: smoking).
Now, the first problem at hand here is the binary classification problem where we aim to differentiate between a smoker and non-smoker using the ECG-based features and other features. For this problem, XGBoost stood out among all four models with an accuracy of 96.5%. The other three models, i.e., Gaussian naive bayes, Decision tree, Random forest achieved relatively low accuracy of 66.2%, 67.4%, 75.2%, respectively. Table IV provides the detailed confusion matrices for all the four classifiers, for this problem.
The next problem at hand is the 4-class classification problem that aims to differentiate between male smokers, female smokers, male non-smokers, and female non-smokers, using the ECG-based features and other features. Among all four models, the multi-layer perceptron performed the best with an accuracy of 97.5%. The decision tree was next in line with an accuracy of 90.5%. The remaining two models, i.e., Gaussian naive bayes, Random forest achieved relatively low accuracy of 66.0%, 87.3%, respectively. Table V provides the detailed confusion matrices for all the four classifiers, for this problem.
Algorithm True Predicted
Non-Smoker
Smoker
Gaussian Naive Bayes
Non-Smoker
Smoker
526
372
42 287
Decision Tree
Non-Smoker
Smoker
363
194
205 465
Random Forest
Non-Smoker
Smoker
374
110
194 549
XGBoost
Non-Smoker
Smoker
541
15
27 644
Table IV: Confusion matrices of the four ML classifiers for the binary classification problem (that aims to differentiate between a smoker and a non-smoker).
Algorithm True Predicted
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
Gaussian Naive Bayes
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
345
89
2
65
42 128
0
67
39 41 130
47
6 15 4 207
Decision Tree
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
832
66
0
75
0 467
0
55
15 5 290
6
6 4 1 631
Random Forest
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
407
98
4
0
0 160
0
4
0 0 123
0
25 15 9 382
Multi Layer Perceptron
Male Smoker
Female Smoker
Male Non-Smoker
Female Non-Smoker
532
10
9
2
8 832
12
0
2 11 742
1
0 0 4 288
Table V: Confusion matrices of the four classifiers for the 4-class classification problem (that aims to differentiate between male smoker, female smoker, male non-smoker, and female non-smoker).
VII Conclusion
This paper presented a novel method to predict the vascular age and smoking status of a person using a custom-built low-cost single-lead ECG sensor module. We constructed our own dataset by collecting ECG data and metadata from 42 apparently healthy subjects (male, female, smoker, non-smoker). We passed the pre-processed and segmented dataset to a number of ML models, a 1D CNN, an MLP, and ResNet18 model. For vascular ageing prediction problem, Random Forest method performed the best with an $R^{2}$ score of 0.99, and MSE of 0.07. For the 2-class smoking classification problem, XGBoost method stood out with an accuracy of 96.5%. Finally, for the 4-class smoking classification problem, MLP method achieved the best accuracy of 97.5%.
This work is aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged (e.g., people in developing countries and people in remote areas). Furthermore, this work opens the door for innovation in wearable devices (e.g., smartwatches and wristbands), enabling accurate and real-time prediction of the vascular age using a single-lead (lead-I) ECG sensor.
As for the future work, we note that our dataset consists of ECG data from apparently healthy young subjects only. Therefore, further efforts are required to collect large amounts of data from a more diverse population (with a much wider age range and with a wide range of cardiovascular diseases). Such dataset once collected, could also help understanding the implications of smoking on cardiovascular health. Further, there is a dire need to validate our results by comparing them with the established classical score-based methods. This investigation will provide valuable insights about the generalization capability of our models in real-world healthcare settings.